OpenStudy (anonymous):

How many different three-digit numbers can you form using the digits 1, 2, 3, 4, 5, 6, and 7 without repetition?

OpenStudy (anonymous):

There are seven possible digits for the first digit and 6 digits for the second (minus one digit for the digit used as the first digit) and 5 options for the last digit (minus one again for the second digit) and then you just multiply them all together to get a total possible combination of 210 numbers that are possible. Read more: http://wiki.answers.com/Q/How_many_different_three_digit_numbers_can_you_form_using_the_digits_1_2_3_5_6_7_and_9_without_repetition#ixzz26lRPeLOg

OpenStudy (anonymous):

7*7 *4 = 49*4 = 196

OpenStudy (anonymous):

thank you